OpenStudy (anonymous):

One integer is 8 more than another. the product is 84. find the integers??? DUE IN TEN MINUTES ONLINE ASSIGNMENT!!!

OpenStudy (anonymous):

6 and 14

OpenStudy (anonymous):

and -14 and -6

OpenStudy (anonymous):

x=8+x... x*(8+x)= 8x+x^2=84.. x^2+8x+84

OpenStudy (anonymous):

x+6& x+14

OpenStudy (anonymous):